OpenStudy (honeymonster):

can you check my answer for finding t in this linear equation..i'll post it in replies

OpenStudy (honeymonster):

\[3\left( t-2 \right)=6\] 3t-6=6 3t-6+6=6+6 3t=12 3t/3=12/3 t=4

OpenStudy (precal):